# Solving General-Utility Markov Decision Processes in the Single-Trial Regime with Online Planning

### v1

@@ 1 Introduction
In this work, we introduce the first approach for solving GUMDPs in the single-trial regime, i.e., when the agent’s performance is evaluated based on a single trial/trajectory. We consider an infinite-horizon discounted setting, which has been greatly adopted by previous research in the field (Zahavy et al., 2021; Hazan et al., 2019) and has found important applications in different domains where the lifetime of the agent is uncertain or infinite. Our key contributions are threefold. First, we establish fundamental results on policy optimization in the single-trial regime, addressing: (i) which class of policies suffices for optimality; (ii) how the problem can be cast as an “occupancy MDP” that is equivalent to our original problem; and (iii) the computational complexity of policy optimization in the single-trial regime. Technically, our results differ from Mutti et al. (2023) due to the inherent differences between infinite-horizon discounted occupancies and the finite-horizon occupancies considered by the previous work. Moreover, our occupancy MDP refines the extended MDP from Mutti et al. (2023), preserves optimality guarantees, and is better suited for practical implementation. Second, we introduce a Monte-Carlo tree search (MCTS) algorithm to solve the occupancy MDP, effectively solving the GUMDP in the single-trial regime via online planning. Our approach provably retrieves the optimal action at each timestep for a sufficiently high number of iterations. Third, we present experimental results showcasing the superior performance of our approach over relevant baselines across diverse tasks (maximum state entropy exploration, imitation learning, and adversarial MDPs) and environments (illustrative GUMDPs and Gym environments (Brockman et al., 2016)).
@@ 2.1 Markov decision processes
MDPs (Puterman, 2014) provide a mathematical framework to study sequential decision making and are formally defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\{{\boldsymbol{P}}^{a}:a\in\mathcal{A}\},{\boldsymbol{p}_{0}},c)$ where: $\mathcal{S}$ is the finite state space; $\mathcal{A}$ is the finite action space; $\{{\boldsymbol{P}}^{a}:a\in\mathcal{A}\}$ is a set of transition probability matrices ${\boldsymbol{P}}^{a}$, one for each action $a\in\mathcal{A}$; ${\boldsymbol{p}_{0}}\in\Delta(\mathcal{S})$ is the initial state distribution; and $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the cost function. For a given action $a\in\mathcal{A}$, each row of matrix ${\boldsymbol{P}}^{a}$ satisfies $P^{a}(s,\cdot)\in\Delta(\mathcal{S})$, encoding the probability of transition from state $s$ at the present timestep to any other state at the next timestep when choosing action $a$. The interaction takes place as follows: (i) an initial state $\mathrm{s}_{0}$ is sampled from ${\boldsymbol{p}}_{0}$; (ii) at each step $t$, the agent observes the state of the environment $\mathrm{s}_{t}\in\mathcal{S}$ and chooses an action $\mathrm{a}_{t}\in\mathcal{A}$. Depending on the chosen action, the environment evolves to state $\mathrm{s}_{t+1}\in\mathcal{S}$ with probability $P^{\mathrm{a}_{t}}(\mathrm{s}_{t},\cdot)$, and the agent receives a random cost $\mathrm{c}_{t}$ with expectation given by $c(\mathrm{s}_{t},\mathrm{a}_{t})$; and (iii) the interaction repeats infinitely.
@@
A policy $\pi=(\pi_{0},\pi_{1},\ldots)$ is a sequence of decision rules, one for each timestep. If, for all timesteps, the decision rules are Markovian or non-Markovian, we say the policy is Markovian or non-Markovian, respectively. Similarly, if the decision rules are deterministic or stochastic for all timesteps, we say the policy is deterministic or stochastic, respectively. We denote the class of non-Markovian policies with $\Pi_{\text{NM}}$, the class of Markovian policies with $\Pi_{\text{M}}$, the class of non-Markovian deterministic policies with $\Pi^{\text{D}}_{\text{NM}}$, and the class of Markovian deterministic policies with $\Pi^{\text{D}}_{\text{M}}$. Finally, the class of stationary policies, $\Pi_{\text{S}}^{\text{D}}$, contains all policies such that the decision rule is the same for all timesteps. We let $\Pi_{\text{S}}$ denote the class of stationary deterministic policies. We highlight that $\Pi_{\text{S}}\subset\Pi_{\text{M}}\subset\Pi_{\text{NM}}$.
For a given policy $\pi\in\Pi_{\text{NM}}$, the interaction between the agent and the environment is a random process $(\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots)$. We let $\mathrm{h}_{t}=(\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots,\mathrm{s}_{t})$ denote a random history up to (including) timestep $t$. We also denote with $h_{t}=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{t})\in\mathcal{S}\times(\mathcal{S}\times\mathcal{A})^{t}$ a particular history up to timestep $t$. The random sequence $(\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots)$ satisfies: (i) $\mathbb{P}\left[\mathrm{s}_{0}=s\right]=p_{0}(s)$; (ii) $\mathbb{P}\left[\mathrm{s}_{t+1}=s^{\prime}|\mathrm{h}_{t},\mathrm{a}_{t}\right]=P^{\mathrm{a}_{t}}(\mathrm{s}_{t},s^{\prime})$; and (iii) $\mathbb{P}\left[\mathrm{a}_{t}=a|\mathrm{h}_{t}\right]=\pi_{t}(a|\mathrm{h}_{t})$. Let $(\Omega,\mathcal{F},\mathbb{P}_{\pi})$ be the probability space over the sequence of random variables $(\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots)$ that satisfies conditions (i)-(iii) above (Lattimore and Szepesvári, 2020). We write specific trajectories as $\omega\in\Omega$, with $\omega=(s_{0},a_{0},s_{1},a_{1},\ldots)$. We denote with $\mathbb{P}_{\pi}\left[\mathrm{s}_{t}=s,\mathrm{a}_{t}=a|\mathrm{s}_{0}\sim{\boldsymbol{p}}_{0}\right]$ the probability of state-action pair $(s,a)$ at timestep $t$ under policy $\pi$.
@@
$d_{\pi}(s,a)$ $=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}_{\pi}\left[\mathrm{s}_{t}=s,\mathrm{a}_{t}=a|\mathrm{s}_{0}\sim{\boldsymbol{p}}_{0}\right].$ (1)
The expected discounted cumulative cost of policy $\pi$ can be written as $J_{\gamma}(\pi)={\boldsymbol{c}}^{\top}{\boldsymbol{d}}_{\pi},$ where ${\boldsymbol{d}}_{\pi}=[d_{\pi}(s_{0},a_{0}),\ldots,d_{\pi}(s_{|\mathcal{S}|},a_{|\mathcal{A}|})]^{\top}$ and ${\boldsymbol{c}}=[c(s_{0},a_{0}),\ldots,c(s_{|\mathcal{S}|},a_{|\mathcal{A}|})]^{\top}.$ Then, the problem of computing the optimal policy becomes $\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi_{\text{S}}}{\boldsymbol{c}}^{\top}{\boldsymbol{d}}_{\pi}$, which can be formulated as a linear program (Puterman, 2014).
@@ 2.2 Monte-Carlo tree search
MCTS (Browne et al., 2012; Silver et al., 2017) is a sample-based planning algorithm to approximate optimal policies in MDPs through sequential tree-based search. The search tree alternates between decision nodes, representing agent actions, and chance nodes, representing stochastic environment transitions. At each iteration, MCTS builds and refines a search tree by alternating between four phases: selection, expansion, simulation, and backpropagation. In the selection phase, the algorithm recursively selects actions at decision nodes according to a tree policy, often based on upper confidence bounds, and samples successor states at chance nodes according to the environment’s dynamics, until it reaches a node that has not yet been fully expanded. Then, in the expansion phase, a new child node corresponding to an unvisited state-action pair is created. In the simulation phase, a rollout policy (typically random or heuristic) generates a trajectory from the expanded node to estimate a Monte Carlo return. Backpropagation then updates the statistics (e.g., mean value, visit counts) along the path traversed during the selection phase. MCTS converges asymptotically to the optimal action at the root under mild assumptions (Kocsis and Szepesvári, 2006).
@@ 2.3 General-utility Markov decision processes
We define an infinite-horizon discounted GUMDP as a tuple $\mathcal{M}_{f}=(\mathcal{S},\mathcal{A},\{{\boldsymbol{P}}^{a}:a\in\mathcal{A}\},{\boldsymbol{p}_{0}},f)$ where $\mathcal{S}$, $\mathcal{A}$, $\{{\boldsymbol{P}}^{a}:a\in\mathcal{A}\}$, and ${\boldsymbol{p}_{0}}$ are defined in a similar way to the standard MDP formulation. The objective of the agent is encoded by $f:\Delta(\mathcal{S}\times\mathcal{A})\rightarrow\mathbb{R}$, as a function of a state-action discounted occupancy ${\boldsymbol{d}}$, as defined in (1). The objective is then to find
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi_{\text{S}}}f({\boldsymbol{d}}_{\pi}).$ (2)
We highlight that, when $f$ is a linear function, we are under the standard MDP setting; if $f$ is convex, then we are under the convex MDP setting (Zahavy et al., 2021). In this work, we consider three different tasks, each associated with a particular (convex) objective function: (i) maximum state entropy exploration (Hazan et al., 2019), where $f({\boldsymbol{d}})={\boldsymbol{d}}^{\top}\log({\boldsymbol{d}})$; (ii) imitation learning (Abbeel and Ng, 2004), where $f({\boldsymbol{d}})=\|{\boldsymbol{d}}-{\boldsymbol{d}}_{\beta}\|_{2}^{2}$ and ${\boldsymbol{d}}_{\beta}\in\Delta(\mathcal{S}\times\mathcal{A})$ is the occupancy induced by behavior policy $\beta$; and (iii) adversarial MDPs (Rosenberg and Mansour, 2019), where $f({\boldsymbol{d}})=\max_{k\in\{1,\ldots,K\}}{\boldsymbol{d}}^{\top}{\boldsymbol{c}}_{k}$ and $\{{\boldsymbol{c}}_{1},\ldots,{\boldsymbol{c}}_{K}\}$ is a set of $K$ cost vectors satisfying $c_{k}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$. Nevertheless, our results apply to any task that can be modelled using the GUMDP framework. We refer to Zahavy et al. (2021) for a comprehensive list of the different objectives considered by previous works.
@@ 2.4 GUMDPs in the single-trial regime
In this work, we consider a different objective from the one introduced in (2). While (2) depends on the expected discounted occupancy, ${\boldsymbol{d}}_{\pi}$, the objective we herein introduce depends on the empirical discounted occupancy induced by running a given policy on the GUMDP. This is particularly important, as practical applications often require identifying the policy that performs optimally when evaluated based on a single trajectory of interaction with the environment. Furthermore, as we shall explain next, in GUMDPs the performance of a given policy may depend on the number of trajectories or trials used to evaluate it (Mutti et al., 2023; Santos et al., 2024).
@@
where $\mathbf{1}$ is the indicator function. In practice, it is common to truncate the trajectories of interaction between the agent and its environment. We denote by $H\in\mathbb{N}$ the truncation horizon and let the empirical truncated occupancy, $\mathbf{d}^{\pi,H}:\Omega\rightarrow\Delta(\mathcal{S}\times\mathcal{A})$, be defined as
@@
$F_{\infty}(\pi)=f({\boldsymbol{d}}_{\pi})=f\left(\mathbb{E}\left[\mathbf{d}^{\pi}\right]\right),$
and we aim to find $\pi^{*}_{\infty}=\operatorname*{arg\,min}_{\pi\in\Pi_{\text{S}}}F_{\infty}(\pi)$. We note that $F_{\infty}$ is equivalent to the objective in (2), which depends on expected occupancies. The fact that $\Pi_{\text{S}}$ suffices for optimality follows from results on the possible state-action occupancies induced by different classes of policies (Puterman, 2014).
@@ 3 Policy Optimization in the Single-Trial Regime
In this section, we establish the fundamental results that underpin the development of online planning algorithms to solve GUMDPs in the single-trial regime. Specifically, we investigate: (i) which class of policies suffices for optimality; (ii) how we can focus on the truncated single-trial objective, $F_{1,H}$, to compute approximately optimal policies for the single trial objective, $F_{1}$; (iii) how we can cast our single-trial GUMDP problem as an MDP in which the agent keeps track of the accrued occupancy at every timestep of the interaction with the GUMDP; and (iv) the computational complexity of policy optimization in the single-trial regime. We let $\mathcal{R}(\pi)=F_{1}(\pi)-\min_{\pi^{\prime}\in\Pi_{\text{NM}}}F_{1}(\pi^{\prime})$ be the regret of an arbitrary policy $\pi\in\Pi_{\text{NM}}$ with respect to the single-trial objective introduced in (5). Intuitively, the regret measures how suboptimal a given policy $\pi$ is compared to the best policy. Throughout our work, we make use of the following assumption.
@@
The objective function $f$ is $L$-Lipschitz with $L\in\mathbb{R}^{+}$, i.e., $|f({\boldsymbol{d}_{1}})-f({\boldsymbol{d}_{2}})|\leq L\|{\boldsymbol{d}}_{1}-{\boldsymbol{d}}_{2}\|_{1}$ for any ${\boldsymbol{d}}_{1},{\boldsymbol{d}}_{2}\in\Delta(\mathcal{S}\times\mathcal{A})$.

### v2

@@ 1 Introduction
In this work, we introduce the first approach for solving GUMDPs in the single-trial regime, i.e., when the agent’s performance is evaluated based on a single trial/trajectory. We consider an infinite-horizon discounted setting, which has been greatly adopted by previous research in the field (Zahavy et al., 2021; Hazan et al., 2019) and has found important applications in different domains where the lifetime of the agent is uncertain or infinite. We focus our attention to environments with discrete state and action spaces. Our key contributions are threefold. First, we establish fundamental results on policy optimization in the single-trial regime, addressing: (i) which class of policies suffices for optimality; (ii) how the problem can be cast as an “occupancy MDP” that is equivalent to our original problem; and (iii) the computational complexity of policy optimization in the single-trial regime. Technically, our results differ from Mutti et al. (2023) due to the inherent differences between infinite-horizon discounted occupancies and the finite-horizon occupancies considered by the previous work. Moreover, our occupancy MDP refines the extended MDP from Mutti et al. (2023), preserves optimality guarantees, and is better suited for practical implementation. Second, we introduce a Monte-Carlo tree search (MCTS) algorithm to solve the occupancy MDP, effectively solving the GUMDP in the single-trial regime via online planning. Our approach provably retrieves the optimal action at each timestep for a sufficiently high number of iterations. Third, we present experimental results showcasing the superior performance of our approach over relevant baselines across diverse tasks and environments.
@@ 2.1 Markov decision processes
MDPs (Puterman, 2014) provide a mathematical framework to study sequential decision making and are formally defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\{{\bm{P}}^{a}:a\in\mathcal{A}\},{\bm{p}_{0}},c)$ where: $\mathcal{S}$ is the finite state space; $\mathcal{A}$ is the finite action space; $\{{\bm{P}}^{a}:a\in\mathcal{A}\}$ is a set of transition probability matrices ${\bm{P}}^{a}$, one for each action $a\in\mathcal{A}$; ${\bm{p}_{0}}\in\Delta(\mathcal{S})$ is the initial state distribution; and $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the cost function. For a given action $a\in\mathcal{A}$, each row of matrix ${\bm{P}}^{a}$ satisfies $P^{a}(s,\cdot)\in\Delta(\mathcal{S})$, encoding the probability of transition from state $s$ at the present timestep to any other state at the next timestep when choosing action $a$. The interaction takes place as follows: (i) an initial state $\mathrm{s}_{0}$ is sampled from ${\bm{p}}_{0}$; (ii) at each step $t$, the agent observes the state of the environment $\mathrm{s}_{t}\in\mathcal{S}$ and chooses an action $\mathrm{a}_{t}\in\mathcal{A}$. Depending on the chosen action, the environment evolves to state $\mathrm{s}_{t+1}\in\mathcal{S}$ with probability $P^{\mathrm{a}_{t}}(\mathrm{s}_{t},\cdot)$, and the agent receives a random cost $\mathrm{c}_{t}$ with expectation given by $c(\mathrm{s}_{t},\mathrm{a}_{t})$; and (iii) the interaction repeats infinitely.
@@
A policy $\pi=(\pi_{0},\pi_{1},\ldots)$ is a sequence of decision rules, one for each timestep. If, for all timesteps, the decision rules are Markovian or non-Markovian, we say the policy is Markovian or non-Markovian, respectively. Similarly, if the decision rules are deterministic or stochastic for all timesteps, we say the policy is deterministic or stochastic, respectively. We denote the class of non-Markovian policies with $\Pi_{\text{NM}}$, the class of Markovian policies with $\Pi_{\text{M}}$, the class of non-Markovian deterministic policies with $\Pi^{\text{D}}_{\text{NM}}$, and the class of Markovian deterministic policies with $\Pi^{\text{D}}_{\text{M}}$. Finally, the class of stationary policies, $\Pi_{\text{S}}$, contains all policies such that the decision rule is the same for all timesteps. We let $\Pi_{\text{S}}^{\text{D}}$ denote the class of stationary deterministic policies.
For a given policy $\pi\in\Pi_{\text{NM}}$, the interaction between the agent and the environment is a random process $(\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots)$. We let $\mathrm{h}_{t}=(\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots,\mathrm{s}_{t})$ denote a random history up to (including) timestep $t$. We also denote with $h_{t}=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{t})\in\mathcal{S}\times(\mathcal{S}\times\mathcal{A})^{t}$ a particular history up to timestep $t$. The random sequence $(\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots)$ satisfies: (i) $\mathbb{P}\left[\mathrm{s}_{0}=s\right]=p_{0}(s)$; (ii) $\mathbb{P}\left[\mathrm{s}_{t+1}=s^{\prime}|\mathrm{h}_{t},\mathrm{a}_{t}\right]=P^{\mathrm{a}_{t}}(\mathrm{s}_{t},s^{\prime})$; and (iii) $\mathbb{P}\left[\mathrm{a}_{t}=a|\mathrm{h}_{t}\right]=\pi_{t}(a|\mathrm{h}_{t})$. Let $(\Omega,\mathcal{F},\mathbb{P}_{\pi})$ be the probability space over the sequence of random variables $(\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots)$ that satisfies conditions (i)-(iii) above (Lattimore and Szepesvári, 2020), where $\mathcal{F}$ is a sigma algebra. We write specific trajectories as $\omega\in\Omega$, with $\omega=(s_{0},a_{0},s_{1},a_{1},\ldots)$. We denote with $\mathbb{P}_{\pi}\left[\mathrm{s}_{t}=s,\mathrm{a}_{t}=a|\mathrm{s}_{0}\sim{\bm{p}}_{0}\right]$ the probability of state-action pair $(s,a)$ at timestep $t$ under policy $\pi$.
@@
$d_{\pi}(s,a)$ $=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}_{\pi}\left[\mathrm{s}_{t}=s,\mathrm{a}_{t}=a|\mathrm{s}_{0}\sim{\bm{p}}_{0}\right].$ (1)
The expected discounted cumulative cost of policy $\pi$ can be written as $J_{\gamma}(\pi)={\bm{c}}^{\top}{\bm{d}}_{\pi},$ where ${\bm{d}}_{\pi}=[d_{\pi}(s_{0},a_{0}),\ldots,d_{\pi}(s_{|\mathcal{S}|},a_{|\mathcal{A}|})]^{\top}$ and ${\bm{c}}=[c(s_{0},a_{0}),\ldots,c(s_{|\mathcal{S}|},a_{|\mathcal{A}|})]^{\top}.$ Then, the problem of computing the optimal policy becomes $\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi_{\text{S}}}{\bm{c}}^{\top}{\bm{d}}_{\pi}$, which can be formulated as a linear program (Puterman, 2014).
@@ 2.2 Monte-Carlo tree search
MCTS (Browne et al., 2012; Silver et al., 2017) is a sample-based planning algorithm to approximate optimal policies in MDPs through sequential tree-based search. The search tree alternates between decision nodes, representing agent actions, and chance nodes, representing stochastic environment transitions. At each iteration, MCTS builds and refines a search tree by alternating between four phases: selection, expansion, simulation, and backpropagation. In the selection phase, the algorithm recursively selects actions at decision nodes according to a tree policy, often based on upper confidence bounds, and samples successor states at chance nodes according to the environment’s dynamics, until it reaches a node that has not yet been fully expanded. Then, in the expansion phase, a new child node corresponding to an unvisited state-action pair is created. In the simulation phase, a rollout policy (typically random or heuristic) generates a trajectory from the expanded node to estimate a Monte Carlo return. Backpropagation then updates the statistics (e.g., mean value, visit counts) along the path traversed during the selection phase. MCTS converges asymptotically to the optimal action at the root under mild assumptions (Kocsis and Szepesvári, 2006). Shah et al. (2020) and Cömer et al. (2025) establish polynomial regret concentration for MCTS-based algorithms.
@@ 2.3 General-utility Markov decision processes
We define an infinite-horizon discounted GUMDP as a tuple $\mathcal{M}_{f}=(\mathcal{S},\mathcal{A},\{{\bm{P}}^{a}:a\in\mathcal{A}\},{\bm{p}_{0}},f)$ where $\mathcal{S}$, $\mathcal{A}$, $\{{\bm{P}}^{a}:a\in\mathcal{A}\}$, and $\bm{p}_{0}$} are defined in a similar way to the standard MDP formulation. The objective of the agent is encoded by $f:\Delta(\mathcal{S}\times\mathcal{A})\rightarrow\mathbb{R}$, as a function of a state-action discounted occupancy $\bm{d}$}, as defined in equation 1. The objective is then to find
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi_{\text{S}}}f({\bm{d}}_{\pi}).$ (2)
We highlight that, when $f$ is a linear function, we are under the standard MDP setting; if $f$ is convex, then we are under the convex MDP setting (Zahavy et al., 2021). In this work, we consider three different tasks, each associated with a particular (convex) objective function: (i) maximum state entropy exploration (Hazan et al., 2019), where $f({\bm{d}})={\bm{d}}^{\top}\log({\bm{d}})$; (ii) imitation learning (Abbeel and Ng, 2004), where $f({\bm{d}})=\|{\bm{d}}-{\bm{d}}_{\beta}\|_{2}^{2}$ and ${\bm{d}}_{\beta}\in\Delta(\mathcal{S}\times\mathcal{A})$ is the occupancy induced by behavior policy $\beta$; and (iii) adversarial MDPs (Rosenberg and Mansour, 2019), where $f({\bm{d}})=\max_{k\in\{1,\ldots,K\}}{\bm{d}}^{\top}{\bm{c}}_{k}$ and $\{{\bm{c}}_{1},\ldots,{\bm{c}}_{K}\}$ is a set of $K$ cost vectors satisfying $c_{k}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$. Nevertheless, our results apply to any task that can be modelled using the GUMDP framework. We refer to Zahavy et al. (2021) for a comprehensive list of the different objectives considered by previous works.
@@ 2.4 GUMDPs in the single-trial regime
In this work, we consider a different objective from the one introduced in equation 2. While equation 2 depends on the expected discounted occupancy, ${\bm{d}}_{\pi}$, the objective we herein introduce depends on the empirical discounted occupancy induced by running a given policy on the GUMDP. This is particularly important, as practical applications often require identifying the policy that performs optimally when evaluated based on a single trajectory of interaction with the environment (Mutti et al., 2023; Santos et al., 2024). Furthermore, as we shall explain next, in GUMDPs the performance of a given policy may depend on the number of trajectories or trials used to evaluate it (Mutti et al., 2023; Santos et al., 2024).
@@
where $\mathbf{1}$ is the indicator function. It holds that $d_{\pi}=\mathbb{E}[\bm{\mathrm{d}}^{\pi}]$, for $d_{\pi}$ as introduced in equation 1. In practice, it is common to truncate the trajectories of interaction between the agent and its environment. We denote by $H\in\mathbb{N}$ the truncation horizon and let the empirical truncated occupancy, $\mathbf{d}^{\pi,H}:\Omega\rightarrow\Delta(\mathcal{S}\times\mathcal{A})$, be defined as
@@
$F_{\infty}(\pi)=f({\bm{d}}_{\pi})=f\left(\mathbb{E}\left[\mathbf{d}^{\pi}\right]\right),$
and we aim to find $\pi^{*}_{\infty}=\operatorname*{arg\,min}_{\pi\in\Pi_{\text{S}}}F_{\infty}(\pi)$. We note that $F_{\infty}$ is equivalent to the objective in equation 2, which depends on expected occupancies. The fact that $\Pi_{\text{S}}$ suffices for optimality follows from results on the possible state-action occupancies induced by different classes of policies (Puterman, 2014).
@@ 3 Policy Optimization in the Single-Trial Regime
In this section, we establish the fundamental results that underpin the development of online planning algorithms to solve GUMDPs in the single-trial regime. Specifically, we investigate: (i) which class of policies suffices for optimality; (ii) how we can focus on the truncated single-trial objective, $F_{1,H}$, to compute approximately optimal policies for the single trial objective, $F_{1}$; (iii) how we can cast our single-trial GUMDP problem as an MDP in which the agent keeps track of the accrued occupancy at every timestep of the interaction with the GUMDP; and (iv) the computational complexity of policy optimization in the single-trial regime. We let $\mathrm{OptGap}(\pi)=F_{1}(\pi)-\min_{\pi^{\prime}\in\Pi_{\text{NM}}}F_{1}(\pi^{\prime})$ be the optimality gap of an arbitrary policy $\pi\in\Pi_{\text{NM}}$ with respect to the single-trial objective introduced in equation 5. Intuitively, the optimality gap measures how suboptimal a given policy $\pi$ is compared to the best policy. Throughout our work, we make use of the following assumption.
@@
The objective function $f$ is $L$-Lipschitz with $L\in\mathbb{R}^{+}$, i.e., $|f({\bm{d}_{1}})-f({\bm{d}_{2}})|\leq L\|{\bm{d}}_{1}-{\bm{d}}_{2}\|_{1}$ for any ${\bm{d}}_{1},{\bm{d}}_{2}\in\Delta(\mathcal{S}\times\mathcal{A})$.